# Gregorio Curello Mannheim Ludvig Sinander

### v1

@@ 1 Introduction
In this paper, we study the comparative statics of dominance: which transformations of payoffs expand or contract the undominated sets? We are inspired by [Weinstein2016, BattigalliEtal2016], who identified sufficient conditions for expansion of the NSD set; we seek conditions that are necessary and sufficient.
@@ 1.1 Related literature
The classic literature on undominatedness (also known as ‘admissibility’) spans statistical decision theory, game theory, and social choice. Much of the focus of this literature is on identifying conditions under which every undominated object is ‘sometimes-optimal’, i.e. expected-payoff-maximising against some (subjective) probability distribution over situations. [^1] See [Wald1939, Stein1955, fishburn1975separation, Pearce1984, Borgers1993, Battigalli1997, BattigalliEtal2016, CheKimKojimaRyan2024, ChengBorgers2025]. We ask a different sort of question, concerning comparative statics.
Two beautiful theorems in [CheKimKojima2021] identify those payoff transformations which lead the NWD set to ‘shift up’ (i.e. become ‘higher’) in a set-wise sense—specifically, in the weak set order. Our comparative-statics question is orthogonal to theirs: what we ask is which transformations of payoffs lead the NWD and NSD sets to expand (or contract).
Most closely related are [Weinstein2016, BattigalliEtal2016], who consider broadly the same comparative-statics question as we do, focusing on strict dominance. [^2] See also [[]section 8.1]deOliveiraLamba2025. What these authors show is that any situation-independent concave and strictly increasing transformation of payoffs expands the NSD set (we recover this result in section˜5.2 below). Our theorem extends and qualifies this finding, by identifying the entire class of payoff transformations which enjoy this expansion property, and also identifying those which enjoy the reverse ‘contraction’ property. We furthermore answer the same questions for the NWD set.
At a higher level, we contribute to the comparative-statics literature [[, e.g.]]Topkis1978,MilgromShannon1994,QuahStrulovici2009. The most closely related paper here is [QuahStrulovici2012], as we detail in a short companion paper [CurelloSinanderWhitmeyer2025note].
@@ 2.1 Comparative risk aversion
In the literature, comparative risk-aversion and its characterisation are almost only ever considered in the special case in which the objects are monetary: $X\subseteq\mathbf{R}$. However, by inspection, properties (a)(c) are equally meaningful whatever the nature of the objects, and in fact they are equivalent even outside of the monetary case, as asserted above. See [CurelloSinanderWhitmeyer2025] for a proof.
@@ 2.2 Sierpiński’s theorem on concave functions
Several of our proofs rely on the following beautiful result due to [Sierpinski1920].
@@
A Lebesgue-measurable function $f:\mathbf{R}\to\mathbf{R}$ is concave if and only if for all $x,y\in\mathbf{R}$, $\frac{1}{2}f(x)+\frac{1}{2}f(y)\leq f(\frac{1}{2}x+\frac{1}{2}y)$.
@@ 3.2 Payoff transformations
Each family $(\phi_{s})_{s\in\mathcal{S}}$ of functions $\mathbf{R}\to\mathbf{R}$ generates a transformation $\Phi:\mathbf{R}^{\mathcal{S}}\to\mathbf{R}^{\mathcal{S}}$ via $(u_{s})_{s\in\mathcal{S}}\mapsto(\phi_{s}(u_{s}))_{s\in\mathcal{S}}$. Such transformations are ‘situation-wise’: for each situation $s\in\mathcal{S}$, the situation-$s$ output $\phi_{s}(u_{s})\in\mathbf{R}$ depends only on the situation-$s$ input $u_{s}\in\mathbf{R}$. Formally, given a family $(\phi_{s})_{s\in\mathcal{S}}$ of functions $\mathbf{R}\to\mathbf{R}$, we say that a transformation $\Phi:\mathbf{R}^{\mathcal{S}}\to\mathbf{R}^{\mathcal{S}}$ is $(\phi_{s})_{s\in\mathcal{S}}$-situation-wise if and only if $\Phi((u_{s})_{s\in\mathcal{S}})=(\phi_{s}(u_{s}))_{s\in\mathcal{S}}$ for every $(u_{s})_{s\in\mathcal{S}}\in\mathbf{R}^{\mathcal{S}}$.
@@ 4 Main theorem
The proofs of the three parts (a)(c) are similar to each other. Sufficiency follows from a short elegant argument substantially due to [Weinstein2016, BattigalliEtal2016] (although these authors use it to establish a weaker claim). Focusing on part (c), this argument goes as follows. The condition in (c) implies, by Pratt’s theorem, that applying $\Phi$ makes the decision-maker less risk-averse in every situation, meaning (by definition) that whenever a mixture $p\in\Delta(X)$ is (strictly) preferred to an object $x\in X$ in a situation $s\in S$ of a finite problem $(X,S,U)$, the same is true in the transformed problem $(X,S,\Phi\circ U)$. Thus any object $x\in X$ which was dominated (by some mixture $p\in\Delta(X)$) in $(X,S,U)$ remains dominated (by that same mixture $p$) in $(X,S,\Phi\circ U)$; in other words, applying $\Phi$ expands the dominated set, so shrinks the undominated set.
@@
By contrast, it takes some work to show that the contractive properties (3) and (4) necessitate situation-wiseness. It is easily shown that these properties necessitate strict situation-wise monotonicity: for any $u,v\in\mathbf{R}^{\mathcal{S}}$ and any situation $s\in\mathcal{S}$, $u_{s}>v_{s}$ implies $\Phi_{s}(u)>\Phi_{s}(v)$. [^8] If $\Phi_{s}(u)\leq\Phi_{s}(v)$ and $u_{s}>v_{s}$, then fixing $x\neq y$ and letting $U(x,\cdot)\coloneqq u$ and $U(y,\cdot)\coloneqq v$ yields $\text{NSD}(\{x,y\},\{s\},U)=\{x\}\nsupseteq\{y\}\subseteq\text{NSD}(\{x,y\},\{s\},\Phi\circ U)$, and similarly NWD. Having shown this, we next argue that for each $s\in\mathcal{S}$, the map $u_{s}\mapsto\Phi_{s}\mathopen{}\mathclose{{\left((u_{t})_{t\in\mathcal{S}}}}\right)$ must be continuous, by showing that any discontinuity would permit the construction of a finite problem $(X,\{s,t\},U)$ that violates properties (3) and (4). Together, strict situation-wise monotonicity and continuity imply situation-wiseness.
@@ 4.1 Situation-wise transformations
The sufficiency (‘if’) claims in Proposition˜1 follow from the ‘short elegant argument’ described immediately after Theorem˜1 above. The necessity of concavity and convexity follows from a kind of reversal of that argument; for part (c), it goes as follows. A failure of convexity means, by Pratt’s theorem, that applying the transformation $\Phi$ fails, in some situation $s\in\mathcal{S}$, to make the decision-maker less risk-averse. By definition, that means that there is a finite problem in which some object $x\in X$ is dis-preferred in situation $s$ to a mixture $p\in\Delta(X)$, but becomes preferred once the transformation is applied. Then (using Sierpiński’s theorem,) a finite problem can be constructed in which $x$ is dominated, but becomes undominated once the transformation is applied—so the undominated set fails to shrink.
Strict monotonicity is necessary in the NWD parts of Proposition˜1 because weak dominance requires ‘weakly better in every situation and strictly better in some’. Hence an object which is not weakly dominated may become weakly dominated when a weakly-but-not-strictly increasing transformation is applied. By contrast, concavity plus weak monotonicity are enough to rule out not-strictly-dominated objects from becoming strictly dominated (because strict dominance demands strict superiority in every situation). Similar logic explains the role of the constant maps.
@@ 5.1 Pareto dominance
Theorem˜1 characterises which payoff transformations generate expansion or contraction of the Pareto frontier $\text{PF}_{S}(X,U)$ of every economy $(X,U)$ for every sub-population $S\subseteq\mathcal{S}$. One upshot is that such transformations must be individual-by-individual, with no scope for one individual’s preferences to influence how another individual’s preferences are transformed. The other upshot (by Pratt’s theorem) is that contraction requires each individual to become less risk-averse, while (neglecting some exceptional cases) expansion requires that each individual become more risk-averse.
@@ 5.2 Strategic dominance in games
We contemplate general transformations $\Phi_{i}:\mathbf{R}^{A}\to\mathbf{R}^{A}$ of a player $i$’s payoff. By Theorem˜1, the transformations $\Phi_{i}$ which contract or expand $\text{NWD}(A_{i},A_{-i},u_{i})$ or $\text{NSD}(A_{i},A_{-i},u_{i})$ in every game $(I,(A_{i},u_{i})_{i\in I})$ must transform player $i$’s payoffs profile-by-profile: for any profile $a_{-i}\in A_{-i}$ of opponents’ actions, player $i$’s transformed payoffs $(\Phi_{i}\circ u_{i})(\cdot,a_{-i}):A_{i}\to\mathbf{R}$ can depend only on the pre-transformation payoffs $u_{i}(\cdot,a_{-i})$ against that same profile $a_{-i}\in A_{-i}$, not on the pre-transformation payoffs $u_{i}(\cdot,b_{-i})$ against any other profile $b_{-i}\in A_{-i}\setminus\{a_{-i}\}$. In terms of the payoff matrix of a two-player game, for the row player, this means a column-by-column transformation, with the transformed entries in each column depending only on the pre-transformed entries in that same column.
@@
The following corollary of Propositions˜1 and 2 gives necessary and sufficient conditions for comparative statics in this special case. The ‘if’ half of part (1) recovers the result of [Weinstein2016, BattigalliEtal2016]; the rest is (presumably) new.

### v2

@@ 1 Introduction
In this paper, we study the comparative statics of dominance: which transformations of payoffs expand or contract the undominated sets? We are inspired by [16, 1], who identified sufficient conditions for expansion of the NSD set; we seek conditions that are necessary and sufficient.
@@ 1.1 Related literature
The classic literature on undominatedness (also known as ‘admissibility’) spans statistical decision theory, game theory, and social choice. Much of the focus of this literature is on identifying conditions under which every undominated object is ‘sometimes-optimal’, i.e. expected-payoff-maximising against some (subjective) probability distribution over situations. [^1] See [15, 14, 9, 11, 3, 2, 1, 5, 6]. We ask a different sort of question, concerning comparative statics.
Two beautiful theorems in [4] identify those payoff transformations which lead the NWD set to ‘shift up’ (i.e. become ‘higher’) in a set-wise sense—specifically, in the weak set order. Our comparative-statics question is orthogonal to theirs: what we ask is which transformations of payoffs lead the NWD and NSD sets to expand (or contract).
Most closely related are [16, 1], who consider broadly the same comparative-statics question as we do, focusing on strict dominance. [^2] See also [[]section 8.1]deOliveiraLamba2025. What these authors show is that any situation-independent concave and strictly increasing transformation of payoffs expands the NSD set (we recover this result in section˜5.2 below). Our theorem extends and qualifies this finding, by identifying the entire class of payoff transformations which enjoy this expansion property, and also identifying those which enjoy the reverse ‘contraction’ property. We furthermore answer the same questions for the NWD set.
At a higher level, we contribute to the comparative-statics literature [[, e.g.]]Topkis1978,MilgromShannon1994,QuahStrulovici2009. The most closely related paper here is [12], as we detail in a separate note [7].
@@ 2.1 Comparative risk aversion
In the literature, comparative risk-aversion and its characterisation are almost only ever considered in the special case in which the objects are monetary: $X\subseteq\mathbf{R}$. However, by inspection, properties (a)(c) are equally meaningful whatever the nature of the objects, and in fact they are equivalent even outside of the monetary case, as asserted above. See [8] for a proof.
@@ 2.2 Sierpiński’s theorem on concave functions
Our proofs rely on the following beautiful result due to [13].
@@
Fix any non-empty open interval $I\subseteq\mathbf{R}$. A Lebesgue-measurable function $f:I\to\mathbf{R}$ is concave if and only if for all $u,v\in I$, $\frac{1}{2}f(u)+\frac{1}{2}f(v)\leq f\bigl(\frac{1}{2}u+\frac{1}{2}v\bigr)$.
@@ 3.2 Payoff transformations
Each family $(\phi_{s})_{s\in\mathcal{S}}$ of functions $\mathbf{R}\to\mathbf{R}$ generates a transformation $\Phi:\mathbf{R}^{\mathcal{S}}\to\mathbf{R}^{\mathcal{S}}$ via $(u_{s})_{s\in\mathcal{S}}\mapsto(\phi_{s}(u_{s}))_{s\in\mathcal{S}}$. Such transformations are ‘situation-wise’: for each situation $s\in\mathcal{S}$, the situation-$s$ output $\phi_{s}(u_{s})\in\mathbf{R}$ depends only on the situation-$s$ input $u_{s}\in\mathbf{R}$. Formally, given a family $(\phi_{s})_{s\in\mathcal{S}}$ of functions $\mathbf{R}\to\mathbf{R}$, we say that a transformation $\Phi:\mathbf{R}^{\mathcal{S}}\to\mathbf{R}^{\mathcal{S}}$ is $(\phi_{s})_{s\in\mathcal{S}}$-situation-wise if and only if $\Phi((u_{s})_{s\in\mathcal{S}})=(\phi_{s}(u_{s}))_{s\in\mathcal{S}}$ for every $(u_{s})_{s\in\mathcal{S}}\in\mathbf{R}^{\mathcal{S}}$. We say that a transformation $\Phi:\mathbf{R}^{\mathcal{S}}\to\mathbf{R}^{\mathcal{S}}$ is situation-wise if and only if it is $(\phi_{s})_{s\in\mathcal{S}}$-situation-wise for some family $(\phi_{s})_{s\in\mathcal{S}}$ of functions $\mathbf{R}\to\mathbf{R}$.
Example 1.
Let $(\phi_{s})_{s\in\mathcal{S}}$ be a family of concave and strictly increasing functions $\mathbf{R}\to\mathbf{R}$, and let $\Phi:\mathbf{R}^{\mathcal{S}}\to\mathbf{R}^{\mathcal{S}}$ be $(\phi_{s})_{s\in\mathcal{S}}$-situation-wise. By Pratt’s theorem, this transformation increases risk-aversion in every situation.
Example 2.
[16, 1] considered the special case of Example˜1 in which risk-aversion is increased in the same way in each situation: there is a single concave and strictly increasing function $\phi:\mathbf{R}\to\mathbf{R}$, and $\Phi:\mathbf{R}^{\mathcal{S}}\to\mathbf{R}^{\mathcal{S}}$ is given by $\Phi((u_{s})_{s\in\mathcal{S}})=(\phi(u_{s}))_{s\in\mathcal{S}}$ for every $(u_{s})_{s\in\mathcal{S}}\in\mathbf{R}^{\mathcal{S}}$.
Example 3.
Let $\phi:\mathbf{R}\to\mathbf{R}$ be strictly increasing but neither concave nor convex (e.g. $u\mapsto u^{3}$), and let $\Phi:\mathbf{R}^{\mathcal{S}}\to\mathbf{R}^{\mathcal{S}}$ be given by $\Phi((u_{s})_{s\in\mathcal{S}})=(\phi(u_{s}))_{s\in\mathcal{S}}$ for every $(u_{s})_{s\in\mathcal{S}}\in\mathbf{R}^{\mathcal{S}}$. By Pratt’s theorem, this transformation neither increases nor decreases risk-aversion.
Example 4.
Fix a probability distribution $\mu\in\Delta(\mathcal{S})$, and let $\Phi:\mathbf{R}^{\mathcal{S}}\to\mathbf{R}^{\mathcal{S}}$ be given by $\Phi((u_{s})_{s\in\mathcal{S}})=\mathopen{}\mathclose{{\left(\sum_{s\in\mathcal{S}}u_{s}\mu(s),\dots,\sum_{s\in\mathcal{S}}u_{s}\mu(s)}}\right)$ for every $(u_{s})_{s\in\mathcal{S}}\in\mathbf{R}^{\mathcal{S}}$. This transformation provides (full) insurance: in any finite problem $(X,S,U)$, each object $x\in X$ now earns $\sum_{s\in\mathcal{S}}U(x,s)\mu(s)$ whatever the situation. Obviously this transformation is not situation-wise.
@@ 4 Main theorem
Theorem˜1 says that situation-wiseness is necessary for a transformation $\Phi$ to robustly expand or robustly contract the undominated sets, that a monotonicity-plus-concavity/convexity property is also necessary for robust expansion/contraction, and that these necessary conditions are jointly sufficient. One implication is that the transformations in Examples˜1 and 2 robustly expand the undominated sets. The transformation in Example˜3 neither robustly expands nor robustly contracts the undominated sets, since it neither increases nor decreases risk-aversion. The insurance-type transformation in Example˜4 also neither robustly expands nor robustly contracts the undominated sets, since it is not situation-wise.
The proofs of the three parts (a)(c) are similar to each other. Sufficiency follows from a short elegant argument substantially due to [16, 1] (although these authors use it to establish a weaker claim). Focusing on part (c), this argument goes as follows. The condition in (c) implies, by Pratt’s theorem, that applying $\Phi$ makes the decision-maker less risk-averse in every situation, meaning (by definition) that whenever a mixture $p\in\Delta(X)$ is (strictly) preferred to an object $x\in X$ in a situation $s\in S$ of a finite problem $(X,S,U)$, the same is true in the transformed problem $(X,S,\Phi\circ U)$. Thus any object $x\in X$ which was dominated (by some mixture $p\in\Delta(X)$) in $(X,S,U)$ remains dominated (by that same mixture $p$) in $(X,S,\Phi\circ U)$; in other words, applying $\Phi$ expands the dominated set, so shrinks the undominated set.
@@
By contrast, it takes some work to show that the contractive properties (3) and (4) necessitate situation-wiseness. It is easily shown that these properties necessitate strict situation-wise monotonicity: for any $u,v\in\mathbf{R}^{\mathcal{S}}$ and any situation $s\in\mathcal{S}$, $u_{s}>v_{s}$ implies $\Phi_{s}(u)>\Phi_{s}(v)$. [^8] If $\Phi_{s}(u)\leq\Phi_{s}(v)$ and $u_{s}>v_{s}$, then fixing $x\neq y$ and letting $U(x,\cdot)\coloneqq v$ and $U(y,\cdot)\coloneqq u$ yields $\text{NSD}(\{x,y\},\{s\},U)\not\ni x\in\text{NSD}(\{x,y\},\{s\},\Phi\circ U)$, and similarly NWD. We then show that for each $s\in\mathcal{S}$, the map $u_{s}\mapsto\Phi_{s}\mathopen{}\mathclose{{\left((u_{t})_{t\in\mathcal{S}}}}\right)$ must be continuous; this and strict situation-wise monotonicity together imply situation-wiseness. To establish continuity, we show that any discontinuity would permit the construction of a finite problem $(\{x,y,z\},\mathcal{S},U)$ that violates properties (3) and (4).
Remark 2.
This continuity argument relies on the richness of the payoff domain. If we were to only contemplate (say) integer payoffs, i.e. if $\Phi$ were a map $\Phi:\mathbf{Z}^{\mathcal{S}}\to\mathbf{Z}^{\mathcal{S}}$, then the continuity argument would fail, and situation-wiseness would in fact not be necessary for properties (3) and (4), as we show in appendix˜H.
@@ 4.1 Situation-wise transformations
The core logic behind Proposition˜1 is most easily understood by focusing on the following special case, in which only situation-independent transformations ($\phi_{s}=\phi$ for every $s\in\mathcal{S}$) and the fully robust properties (1) and (2) are considered.
Corollary 1.
Assume that $2\leq\mathopen{}\mathclose{{\left\lvert\mathcal{S}}}\right\rvert$$\infty$. Fix a function $\phi:\mathbf{R}\to\mathbf{R}$, and let $\Phi:\mathbf{R}^{\mathcal{S}}\to\mathbf{R}^{\mathcal{S}}$ be given by $\Phi((u_{s})_{s\in\mathcal{S}})=(\phi(u_{s}))_{s\in\mathcal{S}}$ for every $(u_{s})_{s\in\mathcal{S}}\in\mathbf{R}^{\mathcal{S}}$.
  • Property (1) holds if and only if $\phi$ is concave and weakly increasing.
  • Property (2) holds if and only if either (i) $\phi$ is concave and strictly increasing or (ii) $\phi$ is constant.
The sufficiency (‘if’) claims in Corollary˜1 follow from the ‘short elegant argument’ described immediately after Theorem˜1 above, except when $\phi$ is constant: in that case, the undominated sets trivially expand when $\Phi$ is applied because $\text{NSD}(X,S,\Phi\circ U)=X=\text{NWD}(X,S,\Phi\circ U)$ for every finite problem $(X,S,U)$.
Proof of necessity (‘only if’) in Corollary˜1.
We first show the necessity of the monotonicity properties, then of concavity. Fix arbitrary $x\neq y\neq z\neq x$. Further fix two situations $s\neq t$ in $\mathcal{S}$ (this is possible since $\mathopen{}\mathclose{{\left\lvert\mathcal{S}}}\right\rvert\geq 2$).
Suppose that $\phi$ is not weakly increasing, i.e. there are $u>v$ such that $\phi(u)<\phi(v)$. Then for $U(x,s)\coloneqq u$ and $U(y,s)\coloneqq v$, it holds that $\text{NSD}(\{x,y\},\{s\},U)\ni x\notin\text{NSD}(\{x,y\},\{s\},\Phi\circ U)$, so property (1) fails.
Suppose that $\phi$ is neither strictly increasing nor constant, i.e. there are $u>v$ and $u^{\prime}\neq v^{\prime}$ such that $\phi(u)\leq\phi(v)$ and $\phi(u^{\prime})<\phi(v^{\prime})$. Then for
[…]
U(row,column) st
we have $\text{NWD}(\{x,y\},\{s,t\},U)\ni x\notin\text{NWD}(\{x,y\},\{s,t\},\Phi\circ U)$, so property (2) fails.
Assume for the remainder that $\phi$ is weakly increasing. Suppose that $\phi$ is not concave. Then by Sierpiński’s theorem from section˜2.2 above (applicable since increasing functions are Lebesgue-measurable), there exist $u>v$ such that $\frac{1}{2}\phi(u)+\frac{1}{2}\phi(v)>\phi\bigl(\frac{1}{2}u+\frac{1}{2}v\bigr)$. Hence for
[…]
U(row,column) st
we have $z\in\text{NWD}(\{x,y,z\},\{s,t\},U)\subseteq\text{NSD}(\{x,y,z\},\{s,t\},U)$, and keeping in mind the mixture $p\in\Delta(\{x,y,z\})$ with $p(x)\coloneqq 1/2\eqqcolon p(y)$, we see that $z\notin\text{NSD}(\{x,y,z\},\{s,t\},\Phi\circ U)\supseteq\text{NWD}(\{x,y,z\},\{s,t\},\Phi\circ U)$. Thus properties (1) and (2) both fail. ∎
@@ 5.1 Pareto dominance
Theorem˜1 characterises which payoff transformations generate expansion or contraction of the Pareto frontier $\text{PF}_{S}(X,U)$ of every economy $(X,U)$ for every sub-population $S\subseteq\mathcal{S}$. One upshot is that such transformations must be individual-by-individual, with no scope for one individual’s preferences to influence how another individual’s preferences are transformed.
This has some striking implications. For example, it is natural to conjecture that increased alignment of interests among individuals in the population will contract the Pareto frontier. Theorem˜1 implies that this conjecture is false, at least when ‘increased alignment’ is formalised as payoffs $U:X\times\mathcal{S}\to\mathbf{R}$ shifting to $\Phi\circ U:X\times\mathcal{S}\to\mathbf{R}$ where $\Phi:\mathbf{R}^{\mathcal{S}}\to\mathbf{R}^{\mathcal{S}}$ is weakly increasing, so that (ceteris paribus) each individual likes an object $x\in X$ more the more it is liked by others. Indeed, situation-wiseness requires that $u_{s}\mapsto\Phi_{t}((u_{r})_{r\in\mathcal{S}})$ be constant for all $s\neq t$ in $\mathcal{S}$, meaning that alignment remains unchanged (neither strictly increases nor strictly decreases).
The other upshot of Theorem˜1 is that (by Pratt’s theorem) contraction requires each individual to become less risk-averse, while (neglecting some exceptional cases) expansion requires that each individual become more risk-averse.
@@ 5.2 Strategic dominance in games
We contemplate general transformations $\Phi_{i}:\mathbf{R}^{A}\to\mathbf{R}^{A}$ of a player $i$’s payoff. By Theorem˜1, the transformations $\Phi_{i}$ which contract or expand $\text{NWD}(A_{i},A_{-i},u_{i})$ or $\text{NSD}(A_{i},A_{-i},u_{i})$ in every game $(I,(A_{i},u_{i})_{i\in I})$ must transform player $i$’s payoffs profile-by-profile: for any profile $a_{-i}\in A_{-i}$ of opponents’ actions, player $i$’s transformed payoffs $(\Phi_{i}\circ u_{i})(\cdot,a_{-i}):A_{i}\to\mathbf{R}$ against a profile $a_{-i}\in A_{-i}$ can depend only on the pre-transformation payoffs $u_{i}(\cdot,a_{-i})$ against that same profile $a_{-i}$, not on the pre-transformation payoffs $u_{i}(\cdot,b_{-i})$ against any other profile $b_{-i}\in A_{-i}\setminus\{a_{-i}\}$. In terms of the payoff matrix of a two-player game, for the row player, this means a column-by-column transformation, with the transformed entries in each column depending only on the pre-transformed entries in that same column.
@@
The following corollary of Propositions˜1 and 2 gives necessary and sufficient conditions for comparative statics in this special case. The ‘if’ half of part (1) recovers the result of [16, 1]; the rest is (presumably) new.